# A characterization of zero entropy loosely Bernoulli flows via FK-pseudometric

### v1

@@ 5. Topologically loosely Kronecker flows
The proof of Theorem 2 is based on results presented above together with the following theorem whose proof can be found in the book of Hasselblatt and Fisher (Theorem 3.3.34 of [FH19]).
Theorem 5.3.
If a continuous flow is uniquely ergodic, then all but countably many time-t maps are uniquely ergodic.
@@
Let $\Phi$ be a continuous flow. Then the following conditions are equivalent:
  1. (1)
    $\Phi$ is topologically loosely Kronecker;
  2. (2)
    $\Phi$ is uniquely ergodic and $(X,\Phi,\mu)$ is loosely Kronecker, where $\{\mu\}~{}=~{}\operatorname{{\mathcal{M}}_{\Phi}^{e}}(X)$.
@@
By the discrete-time result analogous to Theorem 2 (Theorem 5.7 of [GRK22]), the map $\varphi^{t}$ is topologically loosely Kronecker, that is, $\operatorname{\rho_{FK}}(x,y)=0$ for every $x,y\in X$. Using Proposition 3.14 we obtain $\operatorname{\tilde{\rho}_{FK}}(x,y)=0$ for every $x,y\in X$. ∎
@@ 6. Examples and applications
Horocycle flows on surfaces of constant negative curvature are probably the most studied examples of topologically loosely Kronecker flows (see [Fur06], [Mar77], [Mar78], [Rat78], [Rat79] and others). There are some other famous classes, such as strictly ergodic distal flows, that are also known to be quasi-isometric (see [Fur63]). Since the class of loosely Kronecker flows is closed under isometric extensions (see [ORW82]), strictly ergodic distal flows must be loosely Kronecker with respect to its unique ergodic measure.
@@
If $T:X\to X$ is topologically loosely Kronecker, then $\operatorname{\rho_{FK}}(x,y)=0$ for every $x,y\in X$. By Corollary 3.15 we have $\operatorname{\tilde{\rho}_{FK}}((x,r),(y,s))=0$ for every $(x,r),(y,s)\in X_{T}$. ∎
A direct consequence of the fact that special flows over the same automorphism are Kakutani equivalent (Proposition 2.2 in [Kat77]) is the following. We just need to assure that the special flow is indeed a continuous flow.
@@
A natural question is whether dynamical properties persist after time-changes. While some properties such as topological mixing might not be preserved, others as transitivity and minimality are preserved. Being topologically loosely Kronecker is in the later class and as expected from the definition of Kakutani equivalence it remains preserved after continuous time-changes.
@@
Every time-t map of a weakly mixing loosely Kronecker continuous flow is loosely Kronecker.
Since loosely Kronecker transformations are ergodic, another consequence of Theorem 4.6 is the following.
@@
Acknowledgements. The author would like to thank Philipp Kunde and Dominik Kwietniak for the fruitful conversations. The author was supported by NCN Polonez Bis grant no. 2021/43/P/ST1/02885.

### v2

@@ 5. Topologically loosely Kronecker flows
Using results presented above we can restate and prove Theorem 2.
@@
A continuous flow $\Phi$ on $X$ is topologically loosely Kronecker if and only if $\Phi$ is uniquely ergodic and $(X,\Phi,\mu)$ is loosely Kronecker, where $\{\mu\}~{}=~{}\operatorname{{\mathcal{M}}_{\Phi}^{e}}(X)$.
@@
By the discrete-time result analogous to Theorem 2 (Theorem 5.7 in [GRK22]), the map $\varphi^{t}$ is topologically loosely Kronecker, that is, $\operatorname{\rho_{FK}}(x,y)=0$ for every $x,y\in X$. Using Proposition 3.17 we obtain $\operatorname{\tilde{\rho}_{FK}}(x,y)=0$ for every $x,y\in X$. ∎
@@ 6. Examples and applications
Horocycle flows on surfaces of constant negative curvature are probably the most studied examples of topologically loosely Kronecker flows (see [Fur73], [Mar77], [Mar78], [Rat78], [Rat79] and others). There are some other famous classes, such as strictly ergodic distal flows, that are also known to be quasi-isometric (see [Fur63]). Since the class of loosely Kronecker flows is closed under isometric extensions (see Chapter 7 in [ORW82]), strictly ergodic distal flows must be loosely Kronecker with respect to its unique ergodic measure.
@@
If $T:X\to X$ is topologically loosely Kronecker, then $\operatorname{\rho_{FK}}(x,y)=0$ for every $x,y\in X$. By Corollary 3.18 we have $\operatorname{\tilde{\rho}_{FK}}((x,r),(y,s))=0$ for every $(x,r),(y,s)\in~{}X_{T}$. ∎
The following corollary is a direct consequence of Kakutani equivalence of special flows over the same automorphism(Proposition 2.2 in [Kat77]). We just need to assure that the special flow is indeed a continuous flow.
@@
A natural question is whether dynamical properties persist after time-changes. While some properties such as topological mixing might not be preserved, others as transitivity and minimality are preserved. Being topologically loosely Kronecker is in the latter class and as expected from the definition of Kakutani equivalence it remains preserved by continuous time-changes.
@@
Every time-$t$ map for $t\neq 0$ of a weakly mixing loosely Kronecker continuous flow is loosely Kronecker.
We observe that the class of loosely Kronecker flows is not contained in the class of weakly mixing flows. For example, the suspension flow over an irrational rotation on the circle is loosely Kronecker by definition, and it is not weakly mixing (Proposition 3.4.9 in [FH19]). Since loosely Kronecker transformations are ergodic, another consequence of Theorem 4.6 is the following.
@@
Acknowledgements. The author would like to thank Philipp Kunde and Dominik Kwietniak for the fruitful conversations and the reviewers for the valuable comments that improved this work.